# Prospects for PBR detection of KM3-230213A-like events

### v1

@@ Abstract
POEMMA-Balloon with Radio (PBR) is a scaled-down version of the Probe Of Extreme Multi-Messenger Astrophysics (POEMMA) design, optimized to be flown as a payload on one of NASA’s sub-orbital super pressure balloons circling the Earth above the southern oceans for a mission duration of more than 20 days. One of the main science objectives of PBR is to follow up astrophysical event alerts in search of neutrinos with very high energy ($10^{8}\lesssim E_{\nu}/{\rm GeV}\lesssim 10^{10}$). Of particular interest for anticipated PBR observations, the KM3NeT Collaboration has recently reported the detection of the neutrino KM3-230213A with $10^{8.1}\lesssim E_{\nu}/{\rm GeV}\lesssim 10^{8.9}$. Such an unprecedented event is in tension with upper limits on the cosmic neutrino flux from IceCube and the Pierre Auger Observatory: for a diffuse isotropic neutrino flux there is a $3.5\sigma$ tension between KM3NeT and IceCube measurements, and about $2.6\sigma$ if the neutrino flux originates in transient sources. Therefore, if KM3-230213A was not beginner’s luck, it becomes compelling to consider beyond Standard Model (BSM) possibilities which could lead to a signal at KM3NeT-ARCA but not at IceCube/Auger. We calculate the PBR horizon-range sensitivity to probe BSM physics compatible with observation at KM3NeT-ARCA and non-observation at IceCube/Auger. As an illustration, we consider a particular class of BSM physics models which has been described in the literature as a possible explanation of KM3-230213A.
@@ 1 Introduction
POEMMA-Balloon with Radio (PBR) is a stratospheric balloon mission [5] representing the most advanced pathfinder for future space-based missions, such as the Probe of Multi-Messenger Astrophysics (POEMMA) [6]. PBR is designed to fulfill these above requirements as one of its three main scientific objectives: the search for transient astrophysical neutrinos via a Target-of-Opportunity (ToO) approach following multi-messenger event alerts. In addition, its scientific program includes the observation of high-altitude horizontal air-showers (HAHAs) and the inaugural measurement of the fluorescence emissions from extensive air showers (EASs) generated by ultra-high-energy cosmic rays (UHECRs) from a suborbital platform.
@@
As a pathfinder, PBR builds on the design studies of POEMMA and leverages experience from previous balloon missions, including EUSO-SPB1 [7] and EUSO-SPB2 [8]. The payload integrates three primary instruments: a Fluorescence Camera (FC), a Cherenkov Camera (CC), and a dedicated Radio Instrument (RI). The FC, which records the longitudinal profiles of UHECR-induced air showers, is described in detail elsewhere [9]. The CC is incorporated into the hybrid focal surface of a 1.1 m Schmidt telescope with a field of view (FoV) of 36 (vertical) by 24 (horizontal). Together with the RI, which complements the optical measurements, these instruments form the principal detectors for neutrino signatures such as those associated with event KM3-230213A. A NASA-supplied rotator paired with a custom-designed tilt system facilitates precise positioning of both instruments to any location below the horizon, a critical feature for following up on multi-messenger alerts.
@@
The Cherenkov Camera is designed to capture ultra-fast optical pulses, operating with an integration time of only 10 ns. It uses silicon photomultipliers (SiPMs) with a pixel size of 3 mm×$\times$3 mm, yielding an instantaneous FoV of $0.2^{\circ}$ per pixel. The 2048-pixel camera, with a detection wavelength range from 320–900 nm provides a total FoV of $6^{\circ}$ (vertical) by $12^{\circ}$ (horizontal). This configuration is sufficiently wide to monitor an astrophysical event for over 20 minutes without active tracking. The estimated energy threshold is approximately 500 TeV.
@@
The Radio Instrument (RI) consists of two sinusoidal radio antennas with a broadband gain between 50–500 MHz, based on the design of the Low Frequency instrument for the Payload for Ultra-High Energy (PUEO) experiment and optimized for EAS detection [10]. Mounted beneath the telescope, these antennas cover the entire optical FoV with 5dB points of ±$\pm 30^{\circ}$ from boresight, and are externally triggered by the Cherenkov Camera.
PBR is scheduled to fly as a payload on a NASA Super Pressure Balloon (SPB), which will circumnavigate the Southern Ocean for more than 20 days following its launch from Wanaka, New Zealand, in Spring 2027.
@@
In this paper we calculate the PBR horizon-range sensitivity to detect KM3-230213A-like events. The layout is as follows. In Sec. 2 we rexamine the main characteristics of the KM3-230213A event, in Sec. 3 we explore potential interpretations of KM3-230213A that can be investigated with PBR, and in Sec. 4 we evaluate the sensitivity of PBR to such phenomena.
@@ 2 The KM3-230213A event
The highest energy muon event observed by KM3NeT/ARCA during a 284.7 day observation period (using 21 detection strings) recorded muon energy of $120^{+110}_{-60}$ PeV, with a 90% CL interval of 35–380 PeV [4]. The muon’s trajectory was reconstructed to $0.6^{\circ}$ above the horizon and at $259.8^{\circ}$ in azimuth, with a directional uncertainty is $1.5^{\circ}$ (68% CL). This uncertainty is anticipated to improve to $0.12^{\circ}$ once detector positioning is more precicely determined. The event is labeled KM3-230213A, with event coordinate central values of ${\rm RA}=94.3^{\circ}$ and ${\rm DEC}=-7.8^{\circ}$, and MJD$=59988.0533299$. If the event comes from a steady isotropic neutrino flux, its scaled per flavor flux is $E_{\nu}^{2}\Phi(E_{\nu})=5.8^{+10.1}_{-3.7}\times 10^{-8}$ GeV/(cm2s sr) [4], shown by the red data point in fig. 1.
@@
Muons at sea level lose energy in transit to the detector. The KM3NeT collaboration reports a 300 kmwe column depth from sea level to detector for the most likely track trajectory direction of $0.6^{\circ}$ above the horizon. They report that the trajectory column depth from sea level to the detector would be $\sim 60$ kmwe (all water) if its direction is $2.6^{\circ}$ above the horizon. A conservative estimate of muon energy loss from the average energy loss formula
$\Biggl{\langle}\frac{dE_{\mu}}{dX}\Biggr{\rangle}\simeq-bE_{\mu}$ (1)
can be obtained using $b=3.5\times 10^{-6}$cm2/g [15, 16]. For the shorter column depth of $X=60$ kmwe, $E_{\mu}^{f}\simeq E_{\mu}^{i}\exp(-bX)=35$ PeV requires an initial energy of $E_{\mu}^{i}\simeq 4.6\times 10^{16}$ GeV ($4.6\times 10^{10}$ PeV).
@@ 3 Possible neutrino origins of KM3-230213A
Having ruled out the possibility that KM3-230213A was engendered by an atmospheric muon, the next step is to investigate the possible sources of cosmic neutrinos. Actually, a neutrino origin of KM3-230213A is also challenging because the Pierre Auger Observatory and IceCube, which have been operating with a much larger effective area than KM3NeT and for a considerably longer time, have not observed any neutrinos with $E_{\nu}\gtrsim 10^{7.6}~{}{\rm GeV}$ [22, 23]. The absence of extreme energy neutrinos in the Auger [22] or IceCube [23] data samples excludes any association of KM3-230213A with the diffuse cosmogenic neutrino flux. To be more specific, for a diffuse isotropic neutrino flux there is a $3.5\sigma$ tension between KM3NeT-ARCA and IceCube measurements. The tension between KM3NeT-ARCA and IceCube measurements is about $2.6\sigma$ if the neutrino flux originates in a transient source [24, 25]. The combined observations from KM3NeT-ARCA, IceCube, and Auger are consistent with a source flare of duration $T\sim 1~{}{\rm yr}$, with muon neutrino flux $F\sim 2\times 10^{-7}\ (1~{}{\rm yr}/T)~{}{\rm GeV/(cm^{2}s)}$ [26]. The all-sky rate of similar neutrino flaring sources is constrained to be ${\cal R}\lesssim 0.4/{\rm yr}$. Similar results have been reported in [27].
@@
A plethora of models advocating beyond Standard Model (BSM) physics have been proposed to explain the origin of KM3-230213A (we refrain from referring to this literature in detail in these proceedings). Of particular interest for PBR observations are BSM possibilities which could lead to a signal at KM3NeT-ARCA but not at IceCube/Auger. For example, the BSM model conjectured in [35] relies on the fact that the KM3-230213A has traversed approximately 147 km of rock and sea en route to the detector, whereas neutrinos arriving from the same location in the sky would have only traveled through about 14 km of ice before reaching IceCube. This is because the arrival direction of KM3-230213A has an incidence angle of $0.6^{\circ}$ at KM3NeT-ARCA and about $8^{\circ}$ above the horizon at IceCube. The idea proposed in [35] is that a new physics matter potential can resonantly amplify sterile-to-active transitions and accommodate oscillations of extreme energy sterile neutrinos in the 100 km range. The difference in propagation distance can then be used to increase the active neutrino flux near the KM3NeT detector, alleviating the tension between KM3NeT-ARCA and IceCube observations. This BSM model then requires that most of the emission by a point source in the nominal direction of the detected KM3-230213A neutrino arrives at Earth as sterile neutrinos. Two other similar scenarios which make use of the difference in propagation distance are described in [36, 37]. PBR will sample a range of propagation distances. As a point source moves through the FoV of the PBR Cherenkov Camera, particle trajectories through the Earth range up to chord lengths of 2,300 km and column depths up to 7,400 kmwe (see, e.g., fig. 2 in [38]).
In the next section we calculate the horizon-range sensitivity of PBR and show that this stratospheric balloon mission will be able to probe BSM physics compatible with observation at KM3NeT-ARCA and non-observation at IceCube/Auger.
@@ 4 PBR discovery reach
[…]
where $\phi$ and $\delta$ are the co-latitude and longitude of the celestial position of a given source (e.g., $\phi$ is the right ascension in the equatorial celestial coordinate system and $\delta$ is the declination), $T_{0}$ is the observation time, and $f(t)$ is an observation efficiency as a function of time that accounts for the reduction in the observation time due to light from the Sun and Moon [39]. In eq. (2), $A(\beta_{\rm tr}(t,\lambda_{\oplus},\phi_{\oplus}),E_{\nu},\phi,\delta)$ is the instantaneous acceptance to $\nu_{\tau}$s from the source as defined by the geometry of the region of interest around the source on the surface of the Earth and the probability of detecting EASs induced by $\tau$ leptons that from the Earth at a selected location within the region (at e.g., Earth latitude, $\lambda_{\oplus}$, and longitude, ϕsubscriptitalic-$\phi_{\oplus}$) and that propagate along quasi-parallel trajectories that point back to the source. The trajectory of the $\tau$ lepton makes an angle $\beta_{\rm tr}$ above the Earth’s surface. We perform Monte Carlo simulations to sample the region of interest and to calculate the detection probabilities of sampled events based on $\nu_{\tau}$ interaction physics in the Earth, $\tau$-lepton decays, EAS development, and the propagation of Cherenkov light signals through the atmosphere to the detector. In calculating time-averaged acceptances, we account for the changes in the source’s position with respect to the Earth and the detector. We also account for model the balloon’s trajectory based on historical wind patterns and assuming a launch on April 6, 2027.
@@
To determine the PBR discovery reach for neutrino sources, we adopt the quasi-differential neutrino flux over a decade of energy [40], where the sensitivity can be approximated by the relation
${\rm Sensitivity}=\frac{2.44}{\ln 10}\times\frac{N_{\nu}E_{\nu}}{\langle A(E_{\nu})\rangle_{T_{0}}}\ ,$ (3)
where the factor $N_{\nu}=3$ converts the $\nu_{\tau}$ sensitivity to the all-flavor sensitivity, and we have taken the $90\%$ unified confidence level, assuming no signal and no background [41], viz. $2.44/\ln 10$.
Figure 3: All-flavor 90% CL sensitivities for PBR for optimal observing conditions during a 1000-s ToO observation (left) and over a 20-day period limited to dark-sky observations (right). For comparison, the modeled 1000-s and 14-day fluences for a BSM burst are indicated by orange lines and stars. Also plotted are the upper limits for GW170817 reported by IceCube (red), Auger (blue), and ANTARES (magenta) computed for 1000 s and 14 days (see Ref. [42]).
@@
For this flux, we expect a BSM-burst fluence over an observation time of $1000~{}{\rm s}$ to be ${\cal F}_{\rm BSM}\sim 0.1~{}{\rm GeV}\ {\rm cm}^{-2}$ in the energy range of $10^{7.5}\lesssim E_{\nu}/{\rm GeV}\lesssim 10^{9.0}$. For an observation time of 20 days, we expect the fluence to be ${\cal F}_{\rm BSM}\sim 170~{}{\rm GeV}\ {\rm cm}^{-2}$ over this energy range, which is within the projected 90% unified confidence level sensitivity of PBR shown in fig. 3. The population of transient sources proposed in [26] has a cosmological distance scale of 20 Gpc. A precise description of the dependence of the PBR performance in detecting sources at various redshifts will be presented in an upcoming journal publication.

### v3

@@ Abstract
POEMMA-Balloon with Radio (PBR) is a scaled-down version of the Probe Of Extreme Multi-Messenger Astrophysics (POEMMA) design, optimized to be flown as a payload on one of NASA’s sub-orbital super pressure balloons circling the Earth above the southern oceans for a mission duration of more than 20 days. One of the main science objectives of PBR is to follow up astrophysical event alerts in search of neutrinos with very high energy ($10^{8}\lesssim E_{\nu}/{\rm GeV}\lesssim 10^{10}$). Of particular interest for anticipated PBR observations, the KM3NeT Collaboration has recently reported the detection of the neutrino KM3-230213A with $10^{7.8}\lesssim E_{\nu}/{\rm GeV}\lesssim 10^{9.0}$. Such an unprecedented event is in tension with upper limits on the cosmic neutrino flux from IceCube and the Pierre Auger Observatory: for a diffuse isotropic neutrino flux there is a $3.5\sigma$ tension between KM3NeT and IceCube measurements, and about $2.6\sigma$ if the neutrino flux originates in transient sources. Therefore, if KM3-230213A was not beginner’s luck, it becomes compelling to consider beyond Standard Model (BSM) possibilities which could lead to a signal at KM3NeT-ARCA but not at IceCube/Auger. We calculate the PBR horizon-range sensitivity to probe BSM physics compatible with observation at KM3NeT-ARCA and non-observation at IceCube/Auger. As an illustration, we consider a particular class of BSM physics models which has been described in the literature as a possible explanation of KM3-230213A.
@@ 1 Introduction
POEMMA-Balloon with Radio (PBR) is a stratospheric balloon mission [5] representing the most advanced pathfinder for future space-based missions, such as the Probe of Multi-Messenger Astrophysics (POEMMA) [6]. PBR is designed to fulfill these above requirements as one of its three main scientific objectives: the search for transient astrophysical neutrinos via a Target-of-Opportunity (ToO) approach following multi-messenger event alerts. In addition, its scientific program includes the observation of high-altitude horizontal air-showers and the inaugural measurement of the fluorescence emissions from extensive air showers (EASs) generated by ultra-high-energy cosmic rays (UHECRs) from a suborbital platform. PBR is scheduled to fly as a payload on a NASA Super Pressure Balloon (SPB), which will circumnavigate the Southern Ocean for more than 20 days following its launch from Wanaka, New Zealand, in Spring 2027.
@@
As a pathfinder, PBR builds on the design studies of POEMMA and leverages experience from previous balloon missions, including EUSO-SPB1 [7] and EUSO-SPB2 [8]. The payload integrates three primary instruments: a Fluorescence Camera (FC), a Cherenkov Camera (CC), and a dedicated Radio Instrument (RI). The FC, which records the longitudinal profiles of UHECR-induced air showers, is described in detail elsewhere [9]. The CC is incorporated into the hybrid focal surface of a 1.1 m Schmidt telescope with a field of view (FoV) of 36 (vertical) by 24 (horizontal). Together with the RI, which complements the optical measurements, these instruments form the principal detectors for neutrino signatures such as those associated with event KM3-230213A. A NASA-supplied rotator paired with a custom-designed tilt system facilitates precise positioning of both instruments to any location below the horizon.
@@
The CC is designed to capture ultra-fast optical pulses, operating with an integration time of only 10 ns. It uses silicon photomultipliers with a pixel size of $3~{\rm mm}\times 3~{\rm mm}$, yielding an instantaneous FoV of $0.2^{\circ}$ per pixel. The 2048-pixel camera, with a detection wavelength range from 320–900 nm provides a total FoV of $6^{\circ}$ (vertical) by $12^{\circ}$ (horizontal). This configuration is sufficiently wide to monitor an astrophysical event for over 20 minutes without active tracking. The estimated energy threshold is approximately 500 TeV.
@@
The RI consists of two sinusoidal radio antennas with a broadband gain between 50–500 MHz, based on the design of the Low Frequency instrument for the Payload for Ultra-High Energy (PUEO) experiment and optimized for EAS detection [10]. Mounted beneath the telescope, these antennas cover the entire optical FoV with 5dB points of $\pm 30^{\circ}$ from boresight, and are externally triggered by the CC.
@@
In this paper we calculate the PBR sensitivity to KM3-230213A-like events. The layout is as follows. In Sec. 2 we rexamine the main characteristics of the KM3-230213A event, in Sec. 3 we explore potential interpretations of KM3-230213A that can be investigated with PBR, in Sec. 4 we evaluate the sensitivity of PBR to such phenomena, and in Sec. 5 we draw our conclusions.
@@ 2 The KM3-230213A event
The highest energy muon event observed by KM3NeT/ARCA during a 284.7 day observation period (using 21 detection strings) recorded muon energy of $120^{+110}_{-60}$ PeV, with a 90% CL interval of 35–380 PeV [4]. The muon’s trajectory was reconstructed to $0.6^{\circ}$ above the horizon and at $259.8^{\circ}$ in azimuth, with a directional uncertainty is $1.5^{\circ}$ (68% CL). This uncertainty is anticipated to improve to $0.12^{\circ}$ once detector positioning is more precicely determined. The event is labeled KM3-230213A, with event coordinate central values of ${\rm RA}=94.3^{\circ}$ and ${\rm DEC}=-7.8^{\circ}$, and MJD=59988.0533299=59988.0533299. If the event comes from a steady isotropic neutrino flux, its scaled per flavor flux is $E_{\nu}^{2}\Phi(E_{\nu})=5.8^{+10.1}_{-3.7}\times 10^{-8}$ GeV/(cm2s sr) [4], shown by the red data point in Fig. 1.
@@
Muons at sea level lose energy in transit to the detector. The KM3NeT collaboration reports a 300 kmwe column depth from sea level to detector for the most likely track trajectory direction of $0.6^{\circ}$ above the horizon. They report that the trajectory column depth from sea level to the detector would be $\sim 60$ kmwe (all water) if its direction is $2.6^{\circ}$ above the horizon. A conservative estimate of muon energy loss from the average energy loss formula, $\langle dE_{\mu}/dX\rangle\simeq-bE_{\mu}$, can be obtained using $b=3.5\times 10^{-6}$cm2/g [15, 16]. For the shorter column depth of $X=60$ kmwe, $E_{\mu}^{f}\simeq E_{\mu}^{i}\exp(-bX)=35$ PeV requires an initial energy of $E_{\mu}^{i}\simeq 4.6\times 10^{16}$ GeV ($4.6\times 10^{10}$ PeV).
@@ 3 Possible neutrino origins of KM3-230213A
Having ruled out the possibility that KM3-230213A was engendered by an atmospheric muon, the next step is to investigate the possible sources of cosmic neutrinos. Actually, a neutrino origin of KM3-230213A is also challenging because the Pierre Auger Observatory and IceCube, which have been operating with a much larger effective area than KM3NeT and for a considerably longer time, have not observed any neutrinos with $E_{\nu}\gtrsim 10^{7.6}~{\rm GeV}$ [22, 23]. The absence of extreme energy neutrinos in the Auger [22] or IceCube [23] data samples excludes any association of KM3-230213A with the diffuse cosmogenic neutrino flux. To be more specific, for a diffuse isotropic neutrino flux there is a $3.5\sigma$ tension between KM3NeT-ARCA and IceCube measurements. The tension between KM3NeT-ARCA and IceCube measurements is about $2.6\sigma$ if the neutrino flux originates in a transient source [24, 25]. The combined observations from KM3NeT-ARCA, IceCube, and Auger are consistent with a source flare of duration $T\sim 1~{\rm yr}$, with muon neutrino flux $F\sim 2\times 10^{-7}\ (1~{\rm yr}/T)~{\rm GeV/(cm^{2}s)}$ [26]. The all-sky rate of similar neutrino flaring sources is constrained to be ${\cal R}\lesssim 0.4/{\rm yr}$. The population of transient sources proposed in [26] has a cosmological distance scale of 20 Gpc. Similar results have been reported in [27].
@@
A plethora of models advocating beyond Standard Model (BSM) physics have been proposed to explain the origin of KM3-230213A (we refrain from referring to this literature in detail in these proceedings). Of particular interest for PBR observations are BSM possibilities which could lead to a signal at KM3NeT-ARCA but not at IceCube/Auger. For example, the BSM model conjectured in [35] relies on the fact that the KM3-230213A has traversed approximately 147 km of rock and sea en route to the detector, whereas neutrinos arriving from the same location in the sky would have only traveled through about 14 km of ice before reaching IceCube. This is because the arrival direction of KM3-230213A has an incidence angle of $0.6^{\circ}$ at KM3NeT-ARCA and about $8^{\circ}$ above the horizon at IceCube. The idea proposed in [35] is that a new physics matter potential can resonantly amplify sterile-to-active transitions and accommodate oscillations of extreme energy sterile neutrinos in the 100 km range. The difference in propagation distance can then be used to increase the active neutrino flux near the KM3NeT detector, alleviating the tension between KM3NeT-ARCA and IceCube observations. This BSM model then requires that most of the emission by a point source in the nominal direction of the detected KM3-230213A neutrino arrives at Earth as sterile neutrinos. Two other similar scenarios which make use of the difference in propagation distance are described in [36, 37]. PBR will sample a range of propagation distances. As a point source moves through the FoV of PBR’s CC, particle trajectories through the Earth range up to chord lengths of 2,300 km and column depths up to 7,400 kmwe (see, e.g., Fig. 2 in [38]).
@@ 4 PBR discovery reach
$\langle A(E_{\nu},\phi,\delta)\rangle_{T_{0}}=\frac{1}{T_{0}}\int_{t_{0}}^{t_{0}+T_{0}}\,dt\ f(t)\,A(\beta_{\rm tr}(t,\lambda_{\oplus},\phi_{\oplus}),E_{\nu},\phi,\delta)\ ,$ (1)
where $\phi$ and $\delta$ are the co-latitude and longitude of the celestial position of a given source (e.g., $\phi$ is the right ascension in the equatorial celestial coordinate system and $\delta$ is the declination), $T_{0}$ is the observation time, and $f(t)$ is an observation efficiency as a function of time that accounts for the reduction in the observation time due to light from the Sun and Moon [39]. In eq. (1), $A(\beta_{\rm tr}(t,\lambda_{\oplus},\phi_{\oplus}),E_{\nu},\phi,\delta)$ is the instantaneous acceptance to $\nu_{\tau}$s from the source as defined by the geometry of the region of interest around the source on the surface of the Earth and the probability of detecting EASs induced by $\tau$ leptons that from the Earth at a selected location within the region (at e.g., Earth latitude, $\lambda_{\oplus}$, and longitude, $\phi_{\oplus}$) and that propagate along quasi-parallel trajectories that point back to the source. The trajectory of the $\tau$ lepton makes an angle $\beta_{\rm tr}$ above the Earth’s surface. We perform Monte Carlo simulations to sample the region of interest and to calculate the detection probabilities of sampled events based on $\nu_{\tau}$ interaction physics in the Earth, $\tau$-lepton decays, EAS development, and the propagation of Cherenkov light signals through the atmosphere to the detector. In calculating time-averaged acceptances, we account for the changes in the source’s position with respect to the Earth and the detector. We also account for model the balloon’s trajectory based on historical wind patterns and assuming a launch on April 6, 2027.
@@
To determine the PBR discovery reach for neutrino sources, we adopt the quasi-differential neutrino flux over a decade of energy [40]. The $\nu_{\tau}$–sensitivity can be approximated by the relation
$\nu_{\tau}{\rm-sensitivity}=\frac{2.44}{\ln 10}\times\frac{E_{\nu}}{\langle A(E_{\nu})\rangle_{T_{0}}}\ ,$ (2)
where we have taken the $90\%$ CL assuming no signal and no background [41], viz. $2.44/\ln 10$.
Figure 3: Single-flavor 90% CL $\nu_{\tau}$ sensitivity of PBR for optimal observing conditions during a 1000-s ToO observation (left) and over a 20-day period limited to dark-sky observations (right). For comparison, the modeled 1000-s and 14-day fluences for a burst of BSM messengers (which produce $\tau$-leptons inside the Earth) are indicated by orange lines and stars. Also plotted are the upper limits for GW170817 reported by IceCube (red), Auger (blue), and ANTARES (magenta) computed for 1000 s and 14 days (see Ref. [42]).
Figure 4: PBR sky coverage in Galactic coordinates assuming a 20-day flight for a balloon trajectory that follows historical wind patterns. The launch date was set at April 6, 2027 (new moon). The acceptances were calculated at $10^{9}~{\rm GeV}$. For reference, the black stars indicate selected nearby astrophysics sources and reported hotspots from the Telescope Array (TA) experiment, whereas the red star indicates the arrival direction of KM3-230213A.
@@
For this flux of BSM messegers, we expect a fluence over an observation time of $1000~{\rm s}$ to be ${\cal F}_{\rm BSM}\sim 0.1~{\rm GeV}\ {\rm cm}^{-2}$ in the energy range of $10^{7.8}\lesssim E_{\nu}/{\rm GeV}\lesssim 10^{9.0}$. For an observation time of 20 days, we expect the fluence to be ${\cal F}_{\rm BSM}\sim 170~{\rm GeV}\ {\rm cm}^{-2}$ over this energy range. Both fluences are within the projected 90% CL sensitivity of PBR shown in Fig. 3. Furthermore, as one can check by inspection of Fig. 4, PBR’s sky coverage (20 day mission with April 2027 launch) includes the KM3-230213A arrival direction (and a longer flight would cover a larger fraction of the sky).
5 Conclusions
PBR is well-positioned to observe nearby bursting neutrino sources with short duration (viz., emission lasting $\lesssim 1000~{\rm s}$). If KM3-230313A happened to originate from such a nearby transient, with a possible example being the rapid emission of energetic Hawking radiation from a PBH near the end of its evaporation lifetime, then PBR may be able to make a similar observation. On the other hand, KM3NeT and PBR have advantageous chord lengths in the arrival direction of KM3-230213A. This will allow PBR to be sensitive to BSM messengers (which produce $\tau$-leptons upon arrival to Earth) from sources at cosmological distances.